Calculate the power dissipated in the 2 ? resistor, if the voltage of the battarey is ?=27.3 V and the resistance of the resistor is R=5.8 Ohms? Express the answer to two decimal places.2Ω 1Ω R,Ω
Answer #1

find the whole resistance of the circuit
[(5+1)x(2+1)]/[(5+1)+(2+1)]+5.8= 7.8 ohm

2. how much voltage going through the 2 ohm resister
(27.3 V/ 7.8 ohm)x2 ohm = 7V
[7 V/ (2+1)] x 2 ohm = 4.67 V

3. total power dissipate
P=(V`2)/R
P={(4.67)`2}/2
p= 10.90W